Suppose we flip four coins simultaneously: a penny, a nickel, a dime, and a quarter. What is the probability that
(a) they all come up heads?
(b) the penny and nickel both come up heads?
(c) the penny and dime both come up the same?
(d) at least 15 cents wort of coins come up heads?

Answer by stanbon(75887)

Suppose we flip four coins simultaneously: a penny, a nickel, a dime, and a quarter. What is the probability that
(a) they all come up heads?:::: (1/2)^4 = 1/16
--------------------------------------
(b) the penny and nickel both come up heads?
(1/2)(1/2)*1*1 = 1/4
---------------------------------------
(c) the penny and dime both come up the same?
2*(1/4)(1/4)*1*1 = 1/8
---------------------------------------
(d) at least 15 cents worth of coins come up heads?
I'll leave that to you.
Take a look at the 16 points in the sample space
and see how many combinations of coins has value
15 cents or greater.
-----------------------
Cheers,
Stan H.